OpenStudy (anonymous):

If cscθ = 5/4, then cosθ = _____. I thought the answer was 4/5, but I think its wrong

OpenStudy (imstuck):

csc is the reciprocal of sin, not cos.

OpenStudy (imstuck):

csc theta = 5/4, then sin theta = 4/5. Lets look at that:

OpenStudy (imstuck):



OpenStudy (imstuck):

You need to find the value of x before you can find the cos theta. Cos is side adjacent over hypotenuse, and x is the adjacent side. You need to solve that right triangle for x. Can you do that?

OpenStudy (anonymous):

so would it b3 3/4?

OpenStudy (imstuck):

Yep! 3/4 is the cos of that angle. Good!

OpenStudy (anonymous):

awesome!

OpenStudy (imstuck):

No, wait,...it's not 3/4 it's 3/5.

OpenStudy (imstuck):

The hypotenuse is 5 not 4...my oversight mistake. Sorry. x does = 3, but cos is the side adjacent (3) over the hypotenuse (5). 3/5

OpenStudy (anonymous):

oho okay!

OpenStudy (imstuck):

So sorry...hoped to catch you in time!

OpenStudy (anonymous):

you did! thanks!

OpenStudy (imstuck):

You're welcome!